OpenStudy (anonymous):

The population of a type of local dragonfly can be found using an infinite geometric series where a1 = 65 and the common ratio is 1/6. Find the sum of this infinite series that will be the upper limit of this population. 78 28 11 32

OpenStudy (saifoo.khan):

Do you know the formula?

OpenStudy (anonymous):

kind of...

OpenStudy (saifoo.khan):

Formula for..?

OpenStudy (anonymous):

\[\sum_{i=?}^{\infty}.. \] i think that is it..

OpenStudy (anonymous):

this is alg. 2 sigma notation....

OpenStudy (anonymous):

... i dont think that is correct....

OpenStudy (saifoo.khan):

Yes. You can use this formula too.

OpenStudy (anonymous):

... well can u help me work through it?

OpenStudy (saifoo.khan):

Sure.

OpenStudy (saifoo.khan):

So the formula is: \[\Large S_{\infty} = \frac{a}{1-r}\] a is the first term. r is the ratio. Substitute your values.

OpenStudy (anonymous):

\[s \infty=0.16/1-r\]

OpenStudy (saifoo.khan):

a1 = 65 r = 1/6

OpenStudy (anonymous):

my bad, thats what i meant to put.. ive had a long day..

OpenStudy (saifoo.khan):

No problem.

OpenStudy (anonymous):

well can u help me with the answer..? please

OpenStudy (saifoo.khan):

Give it a try.

OpenStudy (anonymous):

i got 78? am i right?

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

Yes. You're correct.